Janny opened a savings account and deposited $300.00 as principal. The account earns 6% interest, compounded annually. What is the balance after 6 years


1
Expert's answer
2021-12-19T14:05:29-0500

"A=P(1+\frac{r}{n})^{nt}"

Compounded annually n=1


"=\$300(1+\frac{0.06}{1})^{1×6}"


"=300(1.06)^6 =\$425.5557"

The balance after 6 years is $425.5557